OpenStudy (anonymous):

Integration question: \[\int\limits_{}^{} {1 \over \sqrt{3x-x^2}}\]

OpenStudy (anonymous):

i've tried allsort of techniques not sure how to go about

OpenStudy (anonymous):

i believe it is linked to arcsin...

OpenStudy (anonymous):

just need a pointer in right direction not full soln

OpenStudy (ash2326):

write 3x-x^2 as \(0-(x^2-3x)\) now complete the square \[(0-(x^2-3x+9/4-9/4)\] \[{(3/2)}^2-{(x-3/2)}^2\] now substitute x-3/2 as u dx = du now use the standard formula

OpenStudy (anonymous):

brilliant

OpenStudy (anonymous):

First of all complete the square in the denominator as : \[\int\limits 1/\sqrt{3x - x^2} dx = \int\limits 1/\sqrt{ [9/4-(x-3/2)^2]} \]

OpenStudy (anonymous):

Now, For the integrand 1/sqrt(9/4-(x-3/2)^2), substitute u = x-3/2 and du = dx

OpenStudy (anonymous):

all right with the rest

OpenStudy (anonymous):

The integral now becomes 1/sqrt(9/4-u^2) du. The integral of 1/sqrt(9/4-u^2) is sin^(-1)((2 u)/3) Now you can proceed easily!